# On action rate admissibility criteria

### v3

@@ 3.2 Isentropic compressible Euler system
To obtain convex integration solutions to (3.1), (3.2) and (3.5), Chiodaroli and Kreml [2] add the fan sub-solution to the maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ of Lemma 3.1. Hence, set $\Lambda=P_{1}$ and take $\tilde{{\boldsymbol{v}}}={\boldsymbol{v}}_{1},\,\tilde{{\boldsymbol{u}}}={\boldsymbol{u}}_{1}$ to obtain
$\partial_{t}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))=0.$ (3.15)
@@
Furthermore, from (3.14) upon elimination of the constant $C$ using Lemma 3.1$(iii)$ we have in $P_{1}$:
$\partial_{t}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))+div_{x}(\rho_{1}({\boldsymbol{u}}+{\boldsymbol{u}}_{1}))+\nabla_{x}\left[p(\rho_{1})+\rho_{1}(({\boldsymbol{v}}+{\boldsymbol{v}}_{1})\otimes({\boldsymbol{v}}+{\boldsymbol{v}}_{1})-({\boldsymbol{u}}+{\boldsymbol{u}}_{1})\right]=0,$ (3.16)
@@
where the trivial term from (3.15) has been retained. The second and last terms in (3.16) cancel and we conclude that $(\rho_{1},\,{\boldsymbol{v}}+{\boldsymbol{v}}_{1})$ satisfy (3.1) and (3.2) in $P_{1}$. An infinite number of weak solutions therefore has been obtained to the Riemann initial value problem with initial conditions (3.5). The corresponding shock speeds depend upon $\bar{\rho}=\rho_{1}$ and are expressed as $\nu_{-}(\rho_{1}),\,\nu_{+}(\rho_{1})$. Explicit expressions are presented in [2, (4.46),(4.47),(4.51),(4.52)]. Moreover, the trace of the expression given in Lemma 3.1 $(iii)$ when $\Lambda=P_{1}$ yields
$C=|\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}}|^{2}\textup{I}_{P_{1}}.$ (3.17)
On the other hand, Chiodaroli and Kreml [2] introduce functions of $\rho_{1}$, $\beta$, $\varepsilon_{1}$, and a parameter $\varepsilon_{2}$ such that $C$ defined in (3.17) can be written as
$C=\beta^{2}(\rho_{1})+\varepsilon_{1}(\rho_{1})+\varepsilon_{2}.$ (3.18)
Let $I$ denote some left interval of $\rho_{m}$ given by
$I=[\rho^{\ast},\,\rho_{m}].$ (3.19)
Then for $\rho_{1}\in I$, the functions $\beta(\rho_{1})$ and $\varepsilon_{1}(\rho_{1})$ are smooth, while $\varepsilon_{2}>0$. In particular,
@@
where $\rho_{m}$ and $v_{m}$ refer respectively to the density and second component of the velocity for the one-dimensional classical 2-shock solution to the Riemann initial value isentropic problem. We have for the 2-shock solution
$\left\{\begin{array}[]{llllll}\rho=\rho_{-},&t>0,&v_{a}=0,&v_{b}=v_{b_{-}},&-\infty<x_{1}<\infty,&x_{2}<\nu_{-}(\rho_{m})t,\\ \rho=\rho_{+},&t>0,&v_{a}=0,&v_{b}=v_{b_{+}},&-\infty<x_{1}<\infty,&x_{2}>\nu_{+}(\rho_{m})t,\\ \rho=\rho_{m},&t>0,&v_{a}=0,&v_{b}=v_{m},&-\infty<x_{1}<\infty,&\nu_{-}(\rho_{m})t<x_{2}<\nu_{+}(\rho_{m})t.\end{array}\right.$
Here, we have set ${\boldsymbol{v}}=(v_{a},\,v_{b})$ and note that $\nu_{-}(\rho_{m}),\,\nu_{+}(\rho_{m}),\,\rho_{-},\,\rho_{+},\,\rho_{m},\,v_{b_{-}},\,v_{b_{+}},,v_{m}$ are determined consistent with the Rankine-Hugoniot conditions.
@@ 4 Justification for LAAP0
In summary, Chiodaroli and Kreml [2] treat the two-dimensional isentropic compressible Euler equation and for specific Riemann initial data prove that there are an infinite number of entropic weak convex integration solutions. For some sub-solutions, the wild solutions constructed in [2] dissipate energy at $t=0^{+}$ more rapidly than the one-dimensional 2-shock solution. As a result. for some sub-solution data the 2-shock solution is not selected by the entropy rate criterion.
In the same Riemann initial value problem, LAAP prefers the 2-shock solution to the entropic wild solutions constructed in [2]. However, Markfelder and Pellhammer [14] exploited Lemma 3.1 to construct examples of solutions which are preferred by LAAP to the 2-shock solution. To this end, note that Lemma 3.1$(i)$ restricts to a compact time interval the two dimensional component of the fan sub-solution, which therefore vanishes smoothly outside this interval. Consequently, for time $T_{0}$ greater than the upper limit of this compact time interval, a new problem may be constructed with piece-wise constant initial data determined by the fan sub-solution. The classical method of elementary waves is then used to extend locally in time the solution obtained in [2]. Hence Markfelder and Pellhammer [14] obtain hybrid solutions whose action for which for sufficiently large finite times is smaller than that of the 2-shock solution. It is worth remarking that the hybrid solution constructed by Markfelder and Pellhammer cannot be trivially extended to a global in time solution as the initial data at $t=T_{0}$ do not necessarily satisfy the small bounded variation hypothesis of Glimm’s global existence theorem [10].
@@ 5 Application of LAAP0
Put
[…]
and note that the 2-shock solution becomes identical to the convex integration solution of [2] in the region external to $\Lambda=P_{1}=[-L_{3},\,L_{3}]\times[\ell_{1},\,\ell_{2}]\times[0,\,t]$ where $2L_{3}$ is defined in Remark 3.1$(ii)$, the left interval $I$ is defined in (3.19),and
$\ell_{1}<\min{(\nu_{-}(\rho_{1}:\rho_{1}\in I)}t,\qquad\ell_{2}>\max{(\nu_{+}(\rho_{1}:\rho_{1}\in I)}t.$ (5.3)
See Figure 1. In consequence, we have
@@
$L_{diff}(\rho_{1})=\frac{1}{2}\rho_{m}v^{2}_{m}-\rho_{m}\varepsilon(\rho_{m})-\frac{1}{2}\rho_{1}C+\rho_{1}\varepsilon(\rho_{1})$ (5.4)
where $C$ is given by (3.18). Set
[…]
@@
[…]
On noting from (5.4) that
$L_{diff}(\rho_{m})=-\frac{1}{2}\rho_{m}\varepsilon_{2}<0,$
we conclude that $D(\rho_{1},t)<0$. Further, because at $t=0$ the domain of integration shown in Figure 1, where the integrand is non-zero, collapses to a set of measure zero, we have
@@
Consider the Riemann initial value problem for the two-dimensional isentropic compressible Euler system (3.1), (3.2) and (3.5). Let the set $S$ of solutions consist of the one-dimensional 2-shock solution, the entropic global convex integration weak solutions constructed by Chiodaroli and Kreml [2], and the hybrid solutions constructed by Markfelder and Pellhammer [14] with fan sub-solution data $\rho_{1}\in I$ and $\varepsilon_{2}>0$. Recall that for small $t>0$ the hybrid solutions coincide with those of [2]. We have
@@
Let $\varepsilon_{2}>0$ and let $\rho_{1}$ belong to a fixed left interval $I$ of $\rho_{m}$ so that a fan sub-solution is defined. Then the 2-shock solution is the sARAC-, and therefore the strictly LAAP0-admissible solution in $S$.
@@
The proof employs components of the proof of [9, Thm. 5], combined with Definitions 2.2 and 2.3. ∎
As just shown, for $\rho_{1}$ in the interval $I$ the action and its first and second time derivatives are lower for the 2-shock solution than for any other solution in $S$, including those obtained in [14] since for small enough time these coincide with those of [2]. Hence, both sARAC and strict LAAP0 exclude the admissibility of the convex integration and hybrid solutions and establish the strict admissibility of the 2-shock solution in $S$.
[…]
@@
Let $\gamma=2.$ For any Riemann data satisfying the necessary condition for the existence of the classical 2-shock solution given by
$(v_{b-}-v_{b+})^{2}\rho_{+}\rho_{-}-(\rho_{+}-\rho_{-})(p(\rho_{+}-p(\rho_{-}))>0,$ (5.7)
and any $\varepsilon_{2}>0$ in (3.18), the classical 2-shock solution is $sARAC$- and strictly LAAP0-admissible in $S$.
@@ Acknowledgements
The authors are grateful for support from the ICMS Research in Groups scheme. They also thank C. M. Dafermos, B. Gebhard and O. Kreml for valuable remarks.

### v4

@@ 3.5 Convex integration solutions
To obtain convex integration solutions to (3.1), (3.2) and (3.5), Chiodaroli and Kreml [2, Proposition 3.1] add the maps $({\boldsymbol{v}},\,{\boldsymbol{u}})$ of Lemma 3.1 to a fan sub-solution. Hence, set $\Lambda=P_{1}$ and take $\tilde{{\boldsymbol{v}}}={\boldsymbol{v}}_{1},\,\tilde{{\boldsymbol{u}}}={\boldsymbol{u}}_{1}$ to obtain from (3.12)
$\partial_{t}\rho_{1}+div_{x}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))=0.$} (3.16)
@@
Furthermore, from (3.14) upon elimination of the constant $C$ using Lemma 3.1(iii)(iii), we have in $P_{1}$:
$\partial_{t}(\rho_{1}({\boldsymbol{v}}+{\boldsymbol{v}}_{1}))+div_{x}(\rho_{1}({\boldsymbol{u}}+{\boldsymbol{u}}_{1}))+{\nabla_{x}p(\rho_{1})+div_{x}\left[\rho_{1}(({\boldsymbol{v}}+{\boldsymbol{v}}_{1})\otimes({\boldsymbol{v}}+{\boldsymbol{v}}_{1})-({\boldsymbol{u}}+{\boldsymbol{u}}_{1})\right]}=0.$ (3.17)
@@
The second and last terms in (3.17) cancel and we conclude that $(\rho_{1},\,{\boldsymbol{v}}+{\boldsymbol{v}}_{1})$ satisfy (3.1) and (3.2) in $P_{1}$ [^1] The derivation given here corrects the one given in our earlier paper [9].. An infinite number of weak solutions therefore has been obtained to the Riemann initial value problem with initial conditions (3.5). We emphasise that this set of weak solutions has been obtained for each fan sub-solution.
3.6 Parameterisation of admissible fan sub-solutions
Description of admissible fan sub-solutions is dealt with in [2, Section 4]. The conclusion we require is that the fan sub-solutions can be parameterised by $\rho_{1}$ and an additional parameter $\varepsilon_{2}\geq 0$; see [2, p. 1035]. Note that $\varepsilon_{2}$ has to satisfy the inequalities [2, (4.82),(4.83)] for admissibility.
We can express the constant $C$ of Lemma 3.1 in terms of these two parameters: the trace of the expression given in Lemma 3.1 $(iii)$ when $\Lambda=P_{1}$ yields
$C=|\tilde{{\boldsymbol{v}}}+{\boldsymbol{v}}|^{2}\qquad\text{{in $P_{1}$.}}$ (3.18)
The constant $C$ in (3.18)) is then given by [2, p. 1043]:
$C=\beta^{2}(\rho_{1})+\varepsilon_{1}(\rho_{1})+\varepsilon_{2},$ (3.19)
where $\beta(\rho_{1})$ is the second component of ${\boldsymbol{v}}_{1}$ and is given by [2, (4.48), (4.53)] and $\varepsilon_{1}(\rho_{1})$ is given by [2, (4.49), (4.50)]. Both $\beta$ and $\varepsilon_{1}$ are smooth functions of $\rho_{1}$,
In the following, $\rho_{m}$ and $v_{m}$ refer, respectively, to the density and second component of the velocity for the intermediate state for the one-dimensional classical 2-shock solution to the Riemann initial value isentropic problem described in Section 3.2. We have (see [2, Lemma 4.5 and p. 1042]):
@@ 3.6 Parameterisation of admissible fan sub-solutions
and $\varepsilon_{2}>0$ on some interval
$I^{\ast}=[\rho^{\ast},\rho_{m}],$ (3.21)
where $\rho^{\ast}$ depends on the Riemann data. We also note that the 2-shock propagates with speeds $\nu_{\pm}(\rho_{m})$; see the discussion below (5.10) in [2].
3.7 Summary
Given the appropriate Riemann initial data, Chiodaroli and Kreml construct admissible fan sub-solutions parameterised by $\rho_{1}$ and $\varepsilon_{2}$. The domain of admissibility of fan sub-solutions is a nonempty domain $G$ in the $(\rho_{1},\varepsilon_{2})$ plane such that $G\subset(\rho_{+},\rho_{m})\times{\mathbb{R}}_{+}$. To each point in $G$ corresponds an uncountable set of convex integration (wild) solutions, which are physically indistinguishable: they all satisfy the same Rankine-Hugoniot conditions and have the same rate of energy dissipation.
Following [2] we denote by $\nu_{-}(\rho_{1}),\nu_{+}(\rho_{1})$ the shock speeds in the $x_{2}$ direction of all wild solutions corresponding to a particular admissible choice of $\rho_{1}$. Explicit expressions are presented in [2, (4.46),(4.47)].
@@ 4 Justification for LAAP0
For some sub-solutions, the wild solutions constructed in [2] dissipate energy at $t=0^{+}$ more rapidly than the one-dimensional 2-shock solution. As a result, for some of their sub-solution data the 2-shock solution is not selected by the entropy rate criterion.
However, in the same Riemann initial value problem, LAAP prefers the 2-shock solution to the entropic wild solutions constructed in [2]. Nevertheless, Markfelder and Pellhammer [14] used Lemma 3.1 to construct solutions which are preferred by LAAP to the 2-shock solution. To this end, they restrict the two dimensional component of the fan sub-solution to a compact time interval. Consequently, for time $T_{0}$ greater than the upper limit of this compact time interval, a new problem may be constructed with piece-wise constant initial data determined by the fan sub-solution. The classical method of elementary waves is then used to extend locally in time the solution obtained in [2]. Hence Markfelder and Pellhammer [14] obtain hybrid solutions whose action for sufficiently large finite times is smaller than that of the 2-shock solution. It is worth remarking that the hybrid solution constructed by Markfelder and Pellhammer cannot be trivially extended to a global in time solution as the initial data at $t=T_{0}$ do not necessarily satisfy the small bounded variation hypothesis of Glimm’s global existence theorem [10].
@@ 5 Application of LAAP0
$A(\rho,\,{\boldsymbol{v}})(t_{0},\,t)=\int^{t}_{t_{0}}\int_{\Omega}\mathcal{L}\,dx\,d\tau,\qquad t>t_{0}.$
The convex integration solution involves the constant $\rho_{1}$ introduced in (3.10) which we now restrict to a sufficiently small left interval $I$ contained in $I^{\ast}$ defined in (3.21). Put
@@
and note that the 2-shock solution becomes identical to the convex integration solution of [2] in the region external to $\Lambda=P_{1}=[-L_{3},\,L_{3}]\times[\ell_{1},\,\ell_{2}]\times[0,\,t]$, where $2L_{3}$ is defined in Remark 3.1(ii)(ii), and
$\ell_{1}<\min{(\nu_{-}(\rho_{1}{)}:\rho_{1}\in I)}t\ {=\nu_{-}(\rho^{\ast})t},\qquad\ell_{2}>\max{(\nu_{+}(\rho_{1}{)}:\rho_{1}\in I)}t\ {=\nu_{+}(\rho^{\ast})t}.$ (5.3)
See Figure 1. The identities involving min, respectively max, here follow from the fact that $\nu_{-}$ is monotone increasing and $\nu_{+}$ is monotone decreasing on $I$, which in turn follows from equations (4.39)-(4.52) of [2].
@@
Set
$D(\rho_{1},\,t)=\int^{L_{3}}_{-L_{3}}\int^{t}_{0}\int^{\ell_{2}}_{\ell_{1}}L_{diff}(\rho_{1})\,dx_{2}d\tau dx_{1}.$
@@
$D(\rho_{m},t)=\int^{L_{3}}_{-L_{3}}\int^{t}_{0}\int^{\nu_{+}(\rho_{m})\tau}_{\nu_{-}(\rho_{m})\tau}L_{diff}(\rho_{m})\,dx_{2}d\tau dx_{1}.$
Observe that in the center wedge in Figure 1 equation (3.18) and (5.1) yield
$L_{diff}(\rho_{1})=\frac{1}{2}\rho_{m}v^{2}_{m}-\rho_{m}\varepsilon(\rho_{m})-\frac{1}{2}\rho_{1}C+\rho_{1}\varepsilon(\rho_{1}),$ (5.4)
where $C$ is given by (3.19). Hence using (3.19) and (3.20) we see that at $\rho_{1}=\rho_{m}$,
$L_{diff}(\rho_{m})=-\frac{1}{2}\rho_{m}\varepsilon_{2}<0.$
We conclude that $D(\rho_{m},t)<0$ for $t>0$. By continuity of $D(\rho_{1},t)$ we have $D(\rho_{1},t)<0$ for $t>0$ when $\rho_{1}$ belongs to a sufficiently small left neighborhood of $\rho_{m}$. Further, because at $t=0$ the domain of integration shown in Figure 1, where the integrand is non-zero, collapses to a set of measure zero, we have
@@
Consider the Riemann initial value problem for the two-dimensional isentropic compressible Euler system (3.1), (3.2) and (3.5). Let the set $S$ of solutions consist of the one-dimensional 2-shock solution, the entropic global convex integration weak solutions constructed by Chiodaroli and Kreml [2], and the hybrid solutions constructed by Markfelder and Pellhammer [14] with fan sub-solution data $\rho_{1}\in I^{\ast}$ and $\varepsilon_{2}>0$. Recall that for small $t>0$ the hybrid solutions coincide with those of [2]. We have
@@
Given Riemann data satisfying (3.6), let $\varepsilon_{2}>0$ and assume that $\rho_{1}$ lies in a sufficiently small left neighborhood $I$ of $\rho_{m}$. Then the 2-shock solution is sARAC-admissible, and therefore is the strictly LAAP0-admissible solution in $S$.
@@
The proof follows by combining the proof of [9, Thm. 5] with Definitions 2.2 and 2.3.
As just shown, for $\rho_{1}$ in the interval $I$ the action and its first and second time derivatives are lower for the 2-shock solution than for any other solution in $S$, including those obtained in [14] since for small enough time these coincide with those of [2]. Hence, both sARAC and strict LAAP0 exclude the admissibility of the convex integration and hybrid solutions and establish the strict admissibility of the 2-shock solution in $S$. ∎
The next result is a global version of Theorem 5.1 corresponding to [9, Thm. 6]. It assumes the pressure law $p(\rho)=\rho^{\gamma}$ with $\gamma=2$ so that the internal energy is $\varepsilon=\rho$. Numerical evidence indicates that an analogous result holds for all $\gamma\in[1,\,3].$
@@
Let $\gamma=2.$ For any Riemann data satisfying (3.6) and any $\varepsilon_{2}>0$ in (3.19), the classical 2-shock solution is sARACsARAC- and strictly LAAP0-admissible in $S$.
@@ Acknowledgements
The authors are grateful for support from the ICMS Research in Groups scheme. They also thank C. M. Dafermos, B. Gebhard and O. Kreml for remarks and the referees for their comments and corrections to an earlier version of this paper.